OpenStudy (curiousquestions):

William is a contractor. He is remodeling his home and has decided to do most of the work himself. Occasionally his family will help with some of the work. William wants to start with replacing the molding on his dining room ceiling and window.

OpenStudy (curiousquestions):

(a) William measures the dining room ceiling to see how much molding he will need to go around the perimeter of the room's ceiling. William measured the length of the dining room to be 8 3/4 ft and the width to be 10 7/8 ft. How much molding does William need for the dining room ceiling? Please show your work.

OpenStudy (anonymous):

You need to add 8 3/4 + 10 7/8

OpenStudy (curiousquestions):

ok.

OpenStudy (curiousquestions):

Woouldn't it be 8 3/4 + 8 3/4 + !0 7/8 +10 7/8 ?

OpenStudy (anonymous):

Yes

OpenStudy (curiousquestions):

ok, thanks

OpenStudy (anonymous):

You can also do it like this \[2 * 8 \frac{3}{4} \space+ \space 2* 10 \frac{7}{8} \] But this works too \[ 8 \frac{3}{4} \space +\space 8 \frac{3}{4} \space +\space 10 \frac{7}{8} \space \space+\space 10 \frac{7}{8}\]

OpenStudy (curiousquestions):

Here's more: (b) William's wife measures the window and adds the measurement to the dining room ceiling measurement. She tells her husband that they need a total of 65 1/8 ft of molding to remold the ceiling and the window. How much molding is needed for the window? Show your work. (c) William begins to plan his work on the window and finds that the pieces for the height of the window will measure 6 3/4 ft each. How long will he cut each piece of molding for the width of the window? Explain your answer. (d) William's wife says the molding for the window will take almost 50% of the total molding they need. William says it is closer to 1/3 of the molding. Approximately what percent of the total molding will be used for the window and who is closer to the correct answer? Explain your answer.

OpenStudy (anonymous):

Note * = times

OpenStudy (curiousquestions):

Great, thanks a lot!

OpenStudy (anonymous):

For B you need to take the answer from a and subtract it by 65 1/8 to find out how much is needed for the window

OpenStudy (curiousquestions):

ok

OpenStudy (curiousquestions):

Here's what I did for (a): (8 3/4 * 2) + (10 7/8 * 2) =17 1/2 + 21 6/8 =17 4/8 + 21 6/8 =39 2/8 =39 1/4 feet

OpenStudy (anonymous):

Yes now 65 1/8 - 39 1/4 = window molding

OpenStudy (curiousquestions):

Here's what I did for (b): 65 1/8 - 39 1/4 =65 1/8 - 39 2/8 =25 7/8 feet Correct?

OpenStudy (anonymous):

Yes

OpenStudy (curiousquestions):

Ok

OpenStudy (anonymous):

Now for c we need to get the width

OpenStudy (curiousquestions):

ok

OpenStudy (anonymous):

25 7/8 - (6 3/4 + 6 3/4) = left over molding once you get that answer you need to divide by two

OpenStudy (curiousquestions):

25 7/8 - (6 3/4 * 2) = width of window Correct?

OpenStudy (anonymous):

make sure to do (6 3/4 + 6 3/4) first and then subtract that from 25 7/8 and then divide by 2

OpenStudy (anonymous):

Once you divide by 2 it will be the width

OpenStudy (curiousquestions):

oh, ok I was close

OpenStudy (curiousquestions):

Yeah, now I understand

OpenStudy (anonymous):

Got it?

OpenStudy (curiousquestions):

I think. Here's what I got for (c): (25 7/8 - (6 3/4 * 2)) / 2 Correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Which should give you 12 3/8

OpenStudy (anonymous):

12 3/8 is the width

OpenStudy (curiousquestions):

Great! Thanks!

OpenStudy (curiousquestions):

I have to show my work step-by-step, so sorry about that.

OpenStudy (anonymous):

It is all good. I am at work so .... :-)

OpenStudy (anonymous):

Was hoping no one called lol

OpenStudy (curiousquestions):

lol

OpenStudy (anonymous):

boss walked by twice but did not say anything and my colleague is talking about lunch lol

OpenStudy (curiousquestions):

XD

OpenStudy (curiousquestions):

Here's what I have so far for (c): (25 7/8 - (6 3/4 * 2 )) / 2 =(25 7/8 - 13 1/2) / 2 =(25 7/8 - 13 4/8) / 2 =12 3/8 / 2 Can you help me with the last part? Also am I correct?

OpenStudy (anonymous):

openstudy is messing up

OpenStudy (curiousquestions):

ok

OpenStudy (anonymous):

Ok lost my self for a minute. 12 3/8 and then divide by 2 like you have it

OpenStudy (anonymous):

You have it?

OpenStudy (anonymous):

Now for d we need to solve who is correct. Is the window 50% (1/2) of the molding or 1/3? total molding = 65 1/8 window molding = 25 7/8

OpenStudy (anonymous):

25 7/8 of 65 1/8 = 25 7/8 * 65 1/8 = decimal / percentage

OpenStudy (anonymous):

ok it is lunch time I hope you get it. Have fun

OpenStudy (curiousquestions):

Thanks a lot!!! Sorry I had to go. Hope to talk to you soon!